Calculate the pH of a buffer made from mixing 7.3 mL of 0.229 M NaC2H3O2 and 12.2 mL of 0.458 M HC2H3O2. The Ka of HC2H3O2 is 1.8 x 10^-5.
Answer #1

Answer:

pH = 4.22

Explanation:

NalgH3 2 First we need to find the number of moles of and HC₂ H₂O₂.Calculation of number of moles of NaC₂H₂O2 Volume = 7.3 mL egiven) Concentration = 0.229 M = 0.229 mol (given) No. of moles

Now we need to calculate pKa from the given value of Ka,

Calculation of pka! Ka= 1.88 10 5 (given) pka = -dog (ka) pka = -log (1.8x103) [pka = 4.74

now, we can use the following equation to calendate the pH of this solution. /No. of males of base) pH = pka + log No. of malthis equation is also known as Henderson Hasselbalch equation.

In this solution, HC2H₂O2 is the acid. Na C₂ H₂O2 is the conjugate base.e No. of moles of base = 1. 6717 mmol No. of moles of acid = 5.587 6 mmol pka = 4.74 Put these values in the equation fore pH